OpenStudy (anonymous):

Given the x-intercept is 4 and y-intercept is also 4, how do you write the equation of the line in the form of ax+by=c ?

hartnn (hartnn):

the 2 intercept form off line equation is , \(\Large \dfrac{x}{a}+\dfrac{y}{b}=1\) where a= x intercept, here a=4 b= y-intercept, here b=4

hartnn (hartnn):

then just multiply throughout by 4 and you'll get it in required form :)

OpenStudy (anonymous):

is the answere \[ \frac{ x }{ a } + \frac{ y }{ b } = 1 ?\]

hartnn (hartnn):

that is the general form, here you are given a=4,b=4 so plug these in that form..

OpenStudy (anonymous):

then, it is \[\frac{ x }{ 4 } + \frac{ y }{ 4 } = 1 ?\]

hartnn (hartnn):

to get the form, ax+by =c you just need to multiply by 4 throughout. this is the last step :)

OpenStudy (anonymous):

where would I multiply 4? Sorry. I have too many questions. I'm not just good at math. Please bare with me.

hartnn (hartnn):

because the form that we require, ax+by=c does not have any denominator. so to eliminate the denominator of 4, we multiply every term by 4

hartnn (hartnn):

so, multiply 4 to each term like this, \(\Large \frac{ x }{ 4 }\times 4 + \frac{ y }{ 4 }\times 4 = 1\times 4 \) can you simplify that ?

OpenStudy (anonymous):

4x + 4y = 4 ?

hartnn (hartnn):

noo...the 4 cancels out, \(\Large \frac{ x }{ \cancel 4 }\times \cancel 4 + \frac{ y }{\cancel 4 }\times \cancel 4 = 1\times 4 \\ \Large \implies x+y=4\) got this ?

OpenStudy (anonymous):

ohh thank you very much. is that also how you do it on other equations?

hartnn (hartnn):

if intercepts are given and we need ax+by=c form, then yes, this is how it is done...

OpenStudy (anonymous):

thank you very much :D it helped me alot.

hartnn (hartnn):

you're welcome ^_^

OpenStudy (anonymous):

Can i ask one last question? What if the intercepts are diff. like -3 and -1 ?

hartnn (hartnn):

yes, then you multiply them first, and multiply the answer throughout. like if you had the intercept, 5 and 7, you would multiply 35 throughout.

OpenStudy (anonymous):

is it like this? \[\frac{ x }{ -3 } \times -3 + \frac{ y }{ -1 } \times -1 = -3\times-1\]

hartnn (hartnn):

no, you should multiply each term by only 1 number. here, you multiplied 1st term by -3 and 2nd term by -1, that is not allowed. let us multiply by -3 throughout, \(\Large \frac{ x }{ -3 } \times -3 + \frac{ y }{ -1 } \times -3 = -3\times1\) now simplify :)

OpenStudy (anonymous):

x + 3y = -1 ?

hartnn (hartnn):

x+3y = -3 -2*1=-3

hartnn (hartnn):

***-3*1=-3

OpenStudy (anonymous):

where did -2 came from?

hartnn (hartnn):

that was typing mistake, i corrected it afterwards., i meant -3*1=-3

OpenStudy (anonymous):

you transposed it?

hartnn (hartnn):

how did u get -1 on right side of x + 3y = -1 ? shouldn't that be -3 ?

OpenStudy (anonymous):

oh my mistake :) yes. it should be -3

hartnn (hartnn):

yes, thats what i was pointing out :) so, x+3y = -3

OpenStudy (anonymous):

oh now i get it. :)) thank you for your patience.

hartnn (hartnn):

you're welcome ^_^

OpenStudy (goformit100):

"A warm Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."